### v1

# Zhening Huang1       Xiaoyang Wu2       Fangcheng Zhong1 Hengshuang Zhao2       Matthias Nießner3       Joan Lasenby1

Abstract
We propose LiteReality, a novel pipeline that converts RGB-D scans of indoor environments into compact, realistic, and interactive 3D virtual replicas. LiteReality not only reconstructs scenes that visually resemble reality but also supports key features essential for graphics pipelines—such as object individuality, articulation, high-quality physically based rendering materials, and physically based interaction. At its core, LiteReality first performs scene understanding and parses the results into a coherent 3D layout and objects, with the help of a structured scene graph. It then reconstructs the scene by retrieving the most visually similar 3D artist-crafted models from a curated asset database. Later, the Material Painting module enhances the realism of retrieved objects by recovering high-quality, spatially varying materials. Finally, the reconstructed scene is integrated into a simulation engine with basic physical properties applied to enable interactive behavior. The resulting scenes are compact, editable, and fully compatible with standard graphics pipelines, making them suitable for applications in AR/VR, gaming, robotics, and digital twins. In addition, LiteReality introduces a training-free object retrieval module that achieves state-of-the-art similarity performance, as benchmarked on the Scan2CAD dataset, along with a robust Material Painting module capable of transferring appearances from images of any style to 3D assets—even in the presence of severe misalignment, occlusion, and poor lighting. We demonstrate the effectiveness of LiteReality on both real-life scans and public datasets.
Figure 1: From RGB-D Scan to Graphics-Ready 3D Scene. LiteReality reconstructs compact, realistic 3D environments from real-world RGB-D scans, featuring near-photorealistic appearance, high-quality articulated geometry, and physically based rendering (PBR) materials—ready for high-fidelity rendering and interactive simulation.
1 Introduction
Creating digital replicas of real-world environments remains a central challenge in machine perception, computer vision, and graphics. Although differentiable rendering techniques have achieved remarkable photorealism in 3D reconstruction mildenhall2020nerf ; kerbl2022efficient , their outputs often resemble “3D photographs”: they capture geometry and texture but lack true interactivity and structure. A truly effective digital replica must go beyond visual fidelity to offer semantically rich, simulation-ready, interactable environments. Such representations empower users to interact with, manipulate, and simulate complex scenes—capabilities essential for robotics, virtual and augmented reality, and embodied AI.
Compared to traditional 3D reconstruction techniques—such as Structure-from-Motion (SfM) schonberger2016structure ; wu2011visualsfm , Gaussian Splatting kerbl2022efficient , and Neural Radiance Fields (NeRF) mildenhall2020nerf ; barron2021mipnerf —we introduce the concept of graphics-ready reconstruction. This form of reconstruction goes beyond visual fidelity by incorporating additional structural and functional information to ensure compatibility with graphics pipelines and support downstream tasks such as interactive simulation. To achieve this, several key criteria must be satisfied. First, the scene should be object-centric, treating each object as an individual entity to enable physically plausible placement and manipulation. Second, object functionality must be modelled, allowing for interactive behaviors such as opening doors, drawers, or appliances. Third, the scene should use PBR materials to ensure photorealistic appearance under varying lighting and relighting conditions. Lastly, the integration of physical properties—including mass, gravity, and collision dynamics—is essential to support realistic interactions grounded in real-world physics. By meeting these requirements, we move beyond passive, geometry-focused reconstructions toward truly graphics-ready environments that combine visual realism with functional structure.
Several research efforts have addressed different aspects of this broader challenge. First, in the simulation domain, recent work study how to create interactive virtual environments from text descriptions Yang_2024_CVPR , room layouts phone2proc , large-scale procedural generation pipelines procthor , and even from single images dai2024acdc . However, these approaches often prioritize diversity over realism, aiming to generate training data for embodied AI rather than faithfully reconstructing digital replicas of real-world scenes. Other methods pursue realism through compact, abstract representations—for instance, using CAD retrieval and alignment to represent scenes izadinia2017im2cad ; gumeli2022roca ; huang2018holistic ; sparc ; mask2cad or estimating object textures and materials from images Yeh_2022_CVPR ; Yan:2023:PSDR-Room ; lopes2024material . While promising, these approaches often overlook the importance of integrating components into a coherent, functioning system and tend to neglect the difficulties posed by real-world indoor scans—such as clutter, severe occlusions, and poor lighting. As a result, existing systems frequently struggle to produce faithful reconstructions from in-the-wild scans, limiting their applicability beyond curated or controlled environments.
To this end, we introduce LiteReality, a method that automatically transforms RGB-D scans into realistic, interactive indoor environments (see Figure 1). LiteReality is designed to handle the diverse challenges of real-life scans. At its core, LiteReality consists of four main stages: 1. Scene Perception and Parsing: Objects and room layouts are extracted from RGB-D scans using off-the-shelf methods and represented in a compact scene graph. This graph enforces spatial constraints and refines noisy input into a structured, physically plausible layout. 2. Object Reconstruction: Artist-designed 3D models that best match real-world objects are retrieved from a curated asset database. Though not perfectly aligned geometrically, these models provide high-fidelity geometry and support interactive functionality, including articulated components. 3. Material Painting: A robust material estimation and optimization framework assigns high-quality PBR materials to the retrieved 3D models, based on multiview RGB images. 4. Procedural Reconstruction: The final scene is assembled by constructing room structures, assembling doors and chairs, placing objects, and assigning basic physical properties and interaction logic—resulting in a lightweight, graphics-ready environment. We evaluate LiteReality on the ScanNet dataset dai2017scannet and on real-world indoor scans captured with an iPhone. Across diverse environments, our method consistently converts noisy, incomplete scans into compact, and realistic digital replicas. Our main contributions are as follows:
  • LiteReality framework: To the best of our knowledge, LiteReality is the first system that converts room-level scans into graphics-ready environments with both high realism and interactive functionality. We demonstrate good generalization across diverse real-world scans.
  • Training-free retrieval pipeline: LiteReality includes a training-free object retrieval system that achieves state-of-the-art (SOTA) similarity performance on the Scan2CAD benchmark.
  • Robust material painting: We introduce a novel material-painting approach that reliably transfers high-fidelity PBR materials onto 3D models—even in the presence of shape misalignments, challenging lighting, and diverse image styles. Quantitative evaluation shows SOTA performance.
2 Related Work
Figure 2: Pipeline of LiteReality. With input RGB-D scans, the process begins with scene perception and parsing, where room layouts and 3D object bounding box are detected and parsed into a structured and physically plausible arrangement with the help of a scene graph. In the object reconstruction stage, identical clustering first finds identical objects, and a hierarchical retrieval approach is conducted to match 3D models retrieved from the LiteReality database. The material painting stage retrieves and optimizes PBR materials by referencing observed images. Finally, procedural reconstruction assembles all the information into a graphics-ready environment that features realism and interactivity.
Graphics-Ready Environments.
While industries such as gaming, AR/VR, film production, and physical simulation rely heavily on traditional computer graphics pipelines to create 3D environments, most state-of-the-art 3D reconstruction methods mildenhall2020nerf ; schonberger2016structure ; kerbl2022efficient are not readily compatible with these pipelines. This incompatibility poses a barrier to their practical adoption in real-world applications. In the research domain, simulation environments designed for embodied AI training are the closest counterparts to graphics-ready environments, as they emphasize both interactivity and realism. Significant efforts have been made to develop such simulation platforms. AI2-THOR ai2thor , RoboTHOR robothor , OmniGibson li2023behavior , Habitat 3.0 puig2023habitat3 , ThreeDWorld threedworld , and VirtualHome puig2018virtualhome are notable examples. Despite their capabilities, these platforms often rely on manually crafted scenes, which limits scalability. Several works have proposed scalable scene generation methods. ProcTHOR procthor extends AI2-THOR with procedurally generated environments; Holodeck Yang_2024_CVPR creates scenes from natural language inputs; Phone2Proc phone2proc generates room layouts by placing semantically similar objects; and Digital Cousin dai2024acdc produces diverse simulated environments from single images. While these approaches improve scalability and diversity, they primarily focus on synthesizing virtual scenes for robotic training rather than faithfully reconstructing real-world environments. This highlights a key gap: building graphics-ready scenes that combine realistic appearance with physical interactivity. A concurrent work, Metascenes yu2025metascenesautomatedreplicacreation , is also working in a similar direction by proposing a large-scale dataset that contains a large number of simulatable scenes constructed from richly annotated ScanNet scans for embodied AI benchmarks. LiteReality, by contrast, is presented as a general-purpose tool for converting real-life scans to graphics-ready scenes—recovering artist-quality PBR materials for realism and generating fully interactive, realistic environments.
Object-Centric Reconstructions.
Joint 3D scene understanding and reconstruction Nie_2020_CVPR ; huang2018cooperative ; lyu2024total ; li2021openrooms is essential for building object-centric scene representations. Recent advances in 3D scene understanding—particularly in oriented bounding box detection lazarow2024cubify ; straub24efm and room layout estimation avetisyan2020scenecad ; li2022deeproom —have made this direction increasingly promising. An emerging line of work focuses on reconstructing object-centric 3D scenes from a single RGB image yao2025castcomponentaligned3dscene ; Yan:2023:PSDR-Room ; dai2024acdc . To model individual objects, some methods perform mesh reconstruction from multi-view RGB images Nie_2020_CVPR ; lyu2024total , while others leverage 3D generative models yao2025castcomponentaligned3dscene . Despite rapid and ongoing progress in these domains, current approaches often suffer from instability or produce low-quality results, limiting their applicability in high-fidelity, simulation-ready applications. An alternative is to retrieve artist-created CAD models, enabling clean, lightweight, and visually coherent scene reconstruction—commonly used in simulation environments. Also, these models often include articulated variants, making them valuable for interactive and setups, though sometimes at the expense of realism. A key direction in this area is CAD model retrieval and alignment for scene reconstruction, as demonstrated by Scan2CAD avetisyan2019scan2cad and many follow-up works izadinia2017im2cad ; gumeli2022roca ; huang2018holistic ; sparc ; mask2cad ; Li2021ODAMOD ; Maninis2020Vid2CADCM ; Tyszkiewicz2022RayTran3P ; Maninis2023CADEstateLC ; langer2024fastcadrealtimecadretrieval ; avetisyan2020scenecad ; Murali2017IndoorSB ; Avetisyan2019EndtoEndCM ; Hampali2021MonteCS ; ainetter2023automatically . Note that these methods primarily emphasize alignment accuracy post-retrieval, often assuming that retrieving any object from the correct category is sufficient. While some works aim to improve retrieval similarity ainetter2023automatically ; Ainetter2024HOCSearch ; Wei2022AccurateInstance , they typically require accurate instance segmentation—a challenging prerequisite in real-world scan data. In contrast, our retrieval method is explicitly designed to address the challenges posed by real-world scans.
Materiel Estimation.
Physically-based rendering (PBR) materials are crucial for achieving photorealism in advanced applications. Estimating PBR materials typically requires recovering spatially varying bidirectional reflectance distribution function (SVBRDF) parameters from RGB images—an inherently ill-posed problem due to the complex interaction of lighting, geometry, and reflectance. A common approach involves retrieval followed by optimization, where materials are initially retrieved based on visual similarity and then refined via differentiable rendering Shi2020:ToG ; li2023end . For instance, PSDR-Room Yan:2023:PSDR-Room and PhotoScene Yeh_2022_CVPR optimize procedural PBR parameters from image observations using differentiable rendering. Similarly, MAPA zhang2024mapa carries out part-level segmentation and style transfer, while Material Palette lopes2024material uses a learning-based framework to estimate plausible materials from segmented object views. Although effective in clearly visible and well-cropped scenarios, these methods struggle in cluttered, room-scale environments. Estimation quality often degrades due to occlusion, inaccurate segmentation, low-resolution inputs, and complex lighting conditions. Additionally, per-object differentiable rendering makes these pipelines computationally expensive, limiting their scalability to scenes with many objects. In LiteReality, we introduce a self-contained and robust method tailored for room-level material estimation in messy, real-world scans. Our method generalizes well under occlusion and partial visibility, and remains reliable across diverse lighting conditions.
3 Methodology
LiteReality is a framework that automatically converts indoor RGB-D scans into Graphics-ready scenes. It begins by performing scene perception to extract the key information of layout and objects. The scene is then procedurally constructed in a physics engine to enable interactivity and simulation capabilities. At its core, the pipeline consists of four key stages: scene perception and parsing, object reconstruction, material painting, and procedural reconstruction, as shown in Figure 2. These stages work together to automatically assemble a complete, graphics-ready environment. We highlight the key features and design of each stage in the following sections.
3.1 Scene Perception and Parsing
Scene Understanding.
LiteReality takes RGB-D images of a room as input. In the scene perception and parsing stage, it extracts key spatial information necessary for producing graphics-ready reconstructions. Among various 3D scene understanding approaches, we identified two key tasks—room layout estimation and oriented bounding box detection—as particularly effective due to their robustness and scalability for generating structured 3D environments. Specifically, for real-life scans, we used Apple’s RoomPlan apple_roomplan_2022 with an iPhone to capture layouts and oriented bounding boxes (O-Bbox) interactively. There are also many open-source alternatives for layout and O-Bbox detection spatiallm ; avetisyan2024scenescript ; straub24efm that can be used. Raw detection results often exhibit noise, such as collisions and misalignments, which can result in issues like overlapping objects, floating items, and discontinuous walls when constructing a 3D world. These problems break physical plausibility and break realism. To address this, we propose a scene graph representation that organizes the scene by explicitly representing spatial and appearance relationships between objects and their surroundings. This structured representation is essential for guiding later constrained collision resolution and ensuring consistency in later stages of the pipeline.
Scene Graph Representation and Parsing.
In the scene graph, nodes represent key elements in the scene, such as walls, windows, doors, and detected objects. Each node is associated with several attributes: 1) spatial properties, including center location $\mathbf{C}\in\mathbb{R}^{3}$ (where $\mathbf{C}=(x,y,z)$), dimensions $\mathbf{D}\in\mathbb{R}^{3}$ (where $\mathbf{D}=(w,h,l)$), and orientation $\boldsymbol{\theta}\in\mathbb{R}^{3}$ (three rotation angles around the $x$, $y$, and $z$ axes); and 2) appearance properties, represented by the top-$k$ most visible cropped images of the object, denoted as a set =$\mathcal{I}=\{I_{1},I_{2},\ldots,I_{k}\}$. These attributes are essential for material estimation, model selection, and layout optimization. Edges between nodes define key spatial relationships and, once constructed, can impose constraints during the spatial arrangement process, ensuring a physically plausible scene. We define four key spatial relationships: 1) attached to walls, for objects physically connected to walls (e.g., shelves, cabinets); 2) on top of, for objects placed on top of others (e.g., a vase on a table); 3) table-chair pair, grouping tables and chairs in typical arrangements; and 4) connecting to, where objects are next to each other and touched. These relationships provide essential constraints for maintaining consistent object placement, preventing collisions, and improving overall scene realism. To compute the edges of the scene graph, the construction process begins by parsing walls into connected components, forming closed polygonal regions that delineate individual rooms. Noisy wall detections are aligned by snapping them to a grid, ensuring well-defined room boundaries. Spatial relationships between objects are inferred based on their center location and orientation, with predefined rules governing each type of relationship. We resolve object collisions iteratively using a Constraint-Based Collision Resolution approach. For each intersecting pair, virtual forces are applied along collision vectors to separate them, while adhering to spatial constraints from the scene graph. For example, wall-attached objects are restricted to wall-aligned movement, and table-chair pairs retain their relative positions. This process continues until collisions are resolved or a maximum iteration count is reached, ensuring a realistic, physically plausible layout. More details can be found in the supplementary materials.
3.2 Object Reconstruction Stage
In the age of ever-expanding 3D asset datasets, it is crucial to develop a robust and scalable retrieval mechanism. Previous approaches langer2024fastcadrealtimecadretrieval ; Avetisyan2019EndtoEndCM ; avetisyan2019scan2cad often rely on pair-wise pre-trained models for retrieval, but these are limited to specific datasets. We believe it is important to adopt a training-free approach, as it would be more adaptable and scalable when the dataset expands. We adopt a hierarchical retrieval approach that leverages multiple information modalities—semantic labels, visual features, and camera pose—to progressively refine the search space and improve accuracy. The retrieval begins with a semantic filtering stage, where objects are categorized by subcategories (e.g., two-seat sofas, bar chairs) to exclude irrelevant candidates. This is followed by an image-based retrieval stage, where cropped 2D views of the input object are compared against pre-rendered images of candidate objects using a pretrained feature encoder (DINOv2oquab2023dinov2 ). The top 10 visually similar matches are selected. In the subsequent pose-aware rendering and comparison step, these candidates are placed into the scene according to the detected pose and rendered from the same camera angles. The resulting views are then cropped and re-encoded for visual feature extraction, further narrowing the selection to the top four candidates. Finally, a contextual selection step employs a language model to assess high-level attributes such as style, proportion, and visual coherence, yielding the best-matched object. This structured, multi-stage pipeline ensures scalable and context-aware retrieval across large datasets. To group identical or highly similar objects, we cluster items within each subcategory using DINOv2 features extracted from cropped images. These features, which capture shape and structure, are averaged per object and clustered via KMeans, with the number of clusters selected by maximizing the silhouette score. Since DINOv2 is color-invariant, we further subdivide clusters by their dominant color to form fine-grained, visually and stylistically consistent groups for joint retrieval.
Figure 3: Procedural Reconstruction Stage. This stage progressively rebuilds rooms by reconstructing layouts, assembling dooors and windows, placing objects, and enabling interactive attributes.
3.3 Material Painting Stage
One of the primary challenges in the pipeline lies in effective and accurate material recovery. While prior approaches Yan:2023:PSDR-Room ; Yeh_2022_CVPR ; lopes2024material have made notable progress using differentiable rendering or prediction-based techniques, they struggle to scale effectively to real-world, noisy environments. This limitation arises from three key factors. First, these methods rely on precise alignment between material segments and cropped image regions. In practice, due to geometric misalignment of objects, their approaches to obtaining reliable crop mappings often fail, resulting in noisy or inconsistent references. Second, poor lighting conditions frequently lead to dark or visually ambiguous crops which, even if correctly mapped to material segments, still result in degraded material initialization. Finally, procedural, graph-based differentiable material rendering introduces significant computational overhead, limiting the ability to process large-scale scenes efficiently. In the Material Painting Stage, we address prior limitations with a compact pipeline that includes automatic crop mapping, multi-step material selection, and lightweight albedo-only optimization. For each retrieved 3D asset, predefined material groups are assigned materials inferred from captured images.
Auto-crop Mapping.
The input to this stage includes multiple cropped image patches and corresponding 3D models. We begin by applying SAM kirillov2023segment -guided segmentation to extract meaningful patches from the images. Specifically, we compute the largest rectangular region within each mask and filter out invalid segments using Grounding DINO detection. Priority is given to smooth, clean regions and larger areas, from which we select the top-k ranked patches. To associate 3D material segments with these image patches, we generate informative visual prompts and employ a Multi-Modal Large Language Model (MLLM) openai2023gpt4 for semantic mapping.
Semantic and Visual Guided Material Search.
To initialize high-quality materials, we combine language-guided and vision-based retrieval from a material database. First, inspired by Make-It-Real fang2024makeitreal , we use multi-step prompting to refine material category predictions and select the top 10 candidates. Then, we extract CLIP embeddings from cropped multi-view patches and compare them with embeddings of the selected materials. Finally, GPT-4 evaluates the visual compatibility of albedo maps with reference patches, ensuring semantic and visual alignment. This multi-stage process produces reliable PBR material initialization for each 3D segment.
Figure 4: Materials Painting Stage Pipeline. Given a 3D model and reference images, this stage recovers realistic PBR materials that enhance visual realism for the 3D object.
Albedo-Only Optimization.
With initial material assignments in place, further refinement is possible using procedural graph-based methods like MaTCH khungurn2015matching or PSDR-Room Yan:2023:PSDR-Room , but these incur high computational costs. Instead, we apply a lightweight, albedo-only adjustment in the CIE LAB color space. Rather than re-optimizing full PBR parameters, we shift the global color distribution of the albedo to better match cropped image patches—preserving high-frequency details while correcting hue and brightness. Formally, let $\mathbf{S}(p)\in\mathbb{R}^{3}$ be the LAB vector of the source albedo at pixel $p$, $\mathbf{T}\in\mathbb{R}^{3}$ the target LAB vector, and $P$ the set of pixels in the albedo map. The adjusted color is:
$\mathbf{S}^{\prime}(p)=\mathbf{S}(p)+\left(\mathbf{T}-\tfrac{1}{|P|}\sum_{q\in P}\mathbf{S}(q)\right).$
This shift aligns the mean color of the albedo with the target while preserving texture details. To infer $\mathbf{T}$, we query an MLLM for part-specific RGB values across views, convert the consensus to LAB, and apply the shift. This efficient post-processing balances visual fidelity and scalability.
3.4 Procedural Reconstruction Stage
Finally, we use a predefined procedural pipeline to reconstruct the scene. This begins with constructing the walls, followed by assembling windows and doors, and finally placing objects in the room. To support physics-based interaction, we assign rigid-body properties to each object, and use their mesh geometry as collision boundaries. Large structural elements, such as walls, are marked as passive rigid bodies, while movable items are set as active rigid bodies. Estimating object mass is crucial for realistic simulation. To achieve this, we crop images of each object and input them into a Multimodal Large Language Model (MLLM), which predicts their mass. These estimates are used to simulate realistic rigid-body dynamics. This entire procedural reconstruction process is conducted within Blender.
Figure 5: Graphics-Ready Reconstruction by LiteReality. Noisy indoor scans are converted into compact, realistic scenes with full PBR materials for all objects, ready for downstream graphics tasks.
4 Experiments and Results
Figure 6: Material Painting Stage Demo: Single reference image applied to multiple 3D models. Close-to-reality PBR materials are recovered, even with sever geometries misalignments.
Figure 7: Material Painting Stage Demo: A 3D model is painted using diverse reference images. Realistic PBR materials are recovered, even under challenging conditions like cartoons or drawings.
We evaluate LiteReality both end-to-end and per stage. Since no prior work directly tackles this task, we propose three benchmarks: retrieval similarity, object-centric PBR material estimation, and full-scene graphics-ready reconstruction. These enable quantitative comparisons with prior work and provide a foundation for future research. We also demonstrate advanced applications of our outputs in the supplementary material.
LiteReality Database.
LiteReality database is structured to align with Apple’s RoomPlan, supporting real-world scene reconstruction across 17 semantic classes. We organized our dataset accordingly, ensuring each category—and relevant subcategories—has a corresponding retrieval database. To build this database, we curated assets primarily from 3D-Future 3d_future and AI2-THOR ai2thor . 3D-Future offers diverse indoor furniture, while AI2-THOR, though smaller, provides detailed articulation data for interactive scenes. For underrepresented categories, we supplemented assets from Sketchfab sketchfab under public-access licenses. As of May 2025, our retrieval database includes 5,283 assets; full statistics are available in the supplementary materials.
Retrieval Similarity Benchmarking.
To evaluate the performance of proposed object retrieval method, we conduct experiments on the ScanNet dataset dai2017scannet . Since Scan2CADavetisyan2019scan2cad provides hand-picked ShapeNetchang2015shapenet models as ground-truth annotations, we perform retrieval over the full ShapeNet database. Following Ainetter2024HOCSearch ; gao2024diffcadweaklysupervisedprobabilisticcad ; Wei2022AccurateInstance , we measure retrieval accuracy using the $L_{1}$ Chamfer Distance between the retrieved model and the ground-truth. Unlike Scan2CAD, which searches within a per-scene CAD pool of roughly 50 candidates, we retrieve from the entire ShapeNet repository, where each category contains 300–3 000 models. Although similar evaluations have been conducted before, direct comparison is hindered by sparse implementation details in prior work. We provide a complete description of our evaluation protocol in the supplementary material to facilitate reproducibility and future comparisons. Experiments use the ScanNet validation scenes.
Object-Centric Material Recovering Benchmarking
BRDF estimation of retrieved objects based on scanned images brings the appearance of real-world objects into the digital space. To evaluate the performance of our proposed method, we established a benchmark using five indoor scenes captured with an iPhone 13 Pro Max running the 3D Scanner App in RoomPlan mode. The scenes include a meeting room, a common area, a bedroom and two study rooms, contains a total of 111 objects across eight categories. For each object, we selected the four most visible frames and tightly cropped each to isolate the object, highlighting real-world scanning challenges such as occlusion, poor lighting, and misalignment. We compare our material painting pipeline against four baselines: (1) PhotoShapephotoshape2018 , a material selection algorithm; (2) Make-It-Realfang2024makeitreal , which use language model to retrieve materials; (3) our Visual+Language Search method; and (4) Make-It-Real with our Albedo-Only Optimization. To enable benchmarking, we also created ground-truth CAD models for the scenes. PBR materials estimated by each method were applied to the models, which are placed at their original poses, and rendered with global illumination using HDR environment maps. Evaluation was based on perceptual differences between the cropped scan images and the corresponding rendered views. More details are presented in the supplementary materials.
Graphics-Ready Reconstruction Benchmarking.
To assess end-to-end reconstruction quality, we introduce a benchmark for graphics-ready reconstruction. Only scenes that are object-centric and realistic (i.e., each object is an individual entity with PBR materials) qualify for graphics-ready reconstruction evaluation. We evaluate four methods in total. First, we reproduce two prior approaches: Phone2Proc phone2proc and Digital Cousin. Second, we extend the Digital Cousin framework by integrating material selection via (1) the original Make-It-Real (MIR) pipeline and (2) our proposed semantic-and-vision guided scheme. All methods are tested on five diverse scenes captured with an iPhone, each containing hundreds of RGB frames. The results are shown in Table 4.
Results.
Table 1: Quantitative Evaluation of Retrieval Similarity. We compute the two-way Chamfer Distance between the normalized GT CAD models annotated in Scan2CAD and our retrieved results. Evaluation is performed on the ScanNet validation set. Our retrieval method consistently outperforms prior approaches, achieving the most accurate shape matches across multiple object categories.
Method avg/CAD ↓ avg/class ↓ bath ↓ bkslf ↓ cab ↓ chr ↓ disp ↓ sfa ↓ tbl ↓ bin ↓
MSCD Wei2022AccurateInstance 0.1103 0.1188 0.1215 0.1114 0.0931 0.1019 0.1423 0.1071 0.1224 0.1119
Digital Cousin dai2024acdc 0.1411 0.1246 0.1439 0.1166 0.1105 0.1363 0.1762 0.1083 0.1832 0.1098
ScanNotate ainetter2023automatically 0.1042 0.1110 0.1161 0.0870 0.0908 0.0995 0.1376 0.1046 0.1183 0.0921
LiteReality 0.0986 0.1067 0.1109 0.0845 0.0859 0.0943 0.1309 0.0951 0.1099 0.0901
Table 2: Object-Centric PBR Material Estimation Comparison. Evaluated on 110 objects across five real-world scanned scenes, LiteReality method demonstrates strong capabilities.
Methods RMSE $\downarrow$ SSIM $\uparrow$ LPIPS $\downarrow$
MIR fang2024makeitreal 0.2377 0.3981 0.6111
PhotoShape photoshape2018 0.3225 0.2371 0.6558
MIR fang2024makeitreal + AO 0.2156 0.4203 0.5899
Sem&Vis 0.2835 0.3758 0.6362
LiteReality 0.2163 0.4353 0.5854
Table 3: Perceptual Quality Evaluation for Full Scenes. Tested by comparing rendered images to captured RGB, LiteReality significantly outperforms prior works across all metrics.
Methods RMSE $\downarrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Phone2Procphone2proc 0.3604 0.5512 0.7338
Digital Cousin dai2024acdc 0.3653 0.5531 0.7364
DC dai2024acdc + Sem&Vis 0.3226 0.5425 0.6717
DC dai2024acdc + MIR fang2024makeitreal 0.3046 0.5492 0.6648
LiteReality 0.2664 0.5818 0.6522
Figure 8: Object-Centric PBR Material Comparison . Note: object pose are unused in MP pipeline
Figure 9: Graphics-Ready Reconstruction Comparison. LiteReality showcases best realism.
As shown in Table 4, our training-free retrieval method achieves the best similarity matching among all baselines. Table 4 quantifies the effectiveness of our material painting approach, clearly outperforming prior methods. For overall scene reconstruction, LiteReality shows strong overall performance in Table 4, achieving the highest visual quality. Figures 5 and 9 illustrate results across diverse real-world scenes and comparisons with baselines. Figure 9 showcases the superior visual results of our proposed methods in PBR comparisons. We further validate robustness in two challenging scenarios: painting multiple 3D models from a single image (Figure 7) and painting a single model using diverse images (Figure 7), including cartoons, drawing or AI-generated content. In both cases, our method produces complete, visually plausible PBR materials that matched that in the image. Additional applications—relighting, editing, physics, AR/VR, and robotics—are shown in the supplementary material.
5 Conclusion
LiteReality demonstrates the ability to reconstruct realistic, interactive indoor scenes from RGB-D scans. Our results showcase the pipeline’s potential in generating visually compelling scenes with physics-based interactions and dynamic lighting. Future work will explore support for additional input modalities, such as single 360-degree images and egocentric videos. Another important direction is deeper physics integration and modelling for dynamic scene interactions and object augmentation. We hope this work serves as a strong foundation for graphics-ready 3D scene reconstruction, with broad applications in robotics, gaming, and AR/VR.
Appendix
Figure 10: Graphics-Ready Reconstruction by LiteReality on ScanNet Dataset Noisy indoor scans are converted into compact, realistic scenes with full PBR materials for all objects, ready for downstream graphics tasks.
Appendix A Outline
In this supplementary document, we present five key sections that provide additional context and depth to our main paper:
  1. 1.
    Database Explanation. We explain the construction of our LiteReality database and discuss how it can be expanded in future work.
  2. 2.
    Benchmark Details. We describe the construction and presentation of three benchmarks, including the rationale behind their design and implementation details to ensure reproducibility.
  3. 3.
    Methodology Details. We provide further description of each stage in our reconstruction pipeline.
  4. 4.
    Additional Results. We present further reconstruction results on the ScanNet dataset, including qualitative examples and accompanying supplementary videos.
  5. 5.
    Application Demonstrations. We showcase several advanced applications enabled by our Graphics-Ready Reconstruction, demonstrating its versatility in the rendering pipeline, virtual reality, robotics, and scene editing.
Finally, we discuss broader limitations of our current system and outline directions for future work.
Appendix B LiteReality Database
The overarching goal in designing LiteReality is to create a robust system capable of converting real-world scans into graphics-ready environments suitable for downstream applications such as VR, robotics, and interactive rendering. Many of our design decisions and technical innovations are driven by the challenge of working with in-the-wild scans. Currently, one of the most reliable solutions for high-quality bounding box and layout detection is Apple’s RoomPlan system apple_roomplan_2022 , which enables interactive indoor scanning with augmented reality support. Our database is thus specifically tailored to be compatible with this capture system.
The RoomPlan framework recognizes 17 key room-defining furniture categories, including bathtub, bed, chair, fireplace, oven, refrigerator, sink, sofa, stairs, storage, stove, table, toilet, and washer-dryer, as well as architectural elements like windows and doors. Accordingly, LiteReality includes these 17 major categories. For broad object types—such as chairs, tables, storage units, and beds—we further divide them into finer subcategories to better reflect the diversity encountered in real-world settings.
Our object assets are sourced from a combination of public datasets, including 3D-FUTURE 3d_future , the articulated dataset from AI2-THOR ai2thor , and several curated handheld scans from Sketchfab sketchfab . For most models, we provide material-level segmentation to support the later Materiel Painting stage. Articulated assets—primarily from AI2-THOR—are carefully processed to maintain sub-mesh structures that accurately represent jointed or movable parts. Note that the number of articulated furniture models remains limited, expanding this subset represents an important direction for future development.
Each 3D model in LiteReality undergoes the following on-boarding steps:
  1. 1.
    Multi-view rendering: We generate object renders from multiple viewpoints and extract DINOv2 features from these images, which are stored for use in later retrieval stages.
  2. 2.
    Orientation normalization: Object orientations are aligned to be front-facing, ensuring consistent placement within axis-aligned bounding boxes.
  3. 3.
    Material segmentation rendering: We extract material segments and generate per-part segmentation renders from multiple viewpoints, then select the most visually representative image for each material segment.
  4. 4.
    Material segmentation visualization: We create informative visualizations of material segmentations by stitching together the representative images for each material part. These visualizations are used in later stages for material matching and retrieval, and have been shown to outperform SOM-based yang2023setofmark labeling methods when used with MLLM prompts (figure 11).
Figure 12 summarizes the database statistics for each subcategory, showcasing material segmentation availability and examples of articulated models with interactive components.
Figure 11: Image stitching for materials image prompts in MLLMs. We tested that image stitching provides more effective visual prompts for MLLMs compared to the commonly used SOM approach.
Figure 12: Literality Database: Statistics and Examples
Appendix C Benchmarking Details
C.1 Retrieval Similarity Evaluation Protocol
Previous work on CAD model retrieval and alignment for 3D scenes—such as Scan2CAD avetisyan2019scan2cad , Vid2CAD Maninis2021Vid2CAD , and FastCAD Langer2024FastCAD —has primarily focused on alignment accuracy. These methods typically follow the Scan2CAD evaluation protocol, which considers a retrieval successful if the retrieved CAD model belongs to the correct object category, and then evaluates performance based on alignment precision.
In contrast, our pipeline does not perform explicit alignment. Instead, we rely on an off-the-shelf detector to predict 3D oriented bounding boxes for objects, making the Scan2CAD alignment-based protocol inapplicable to our setting.
To evaluate retrieval similarity, we adopt more rigorous and fine-grained metrics proposed in recent literature. Works such as DiffCAD gao2024diffcadweaklysupervisedprobabilisticcad , HOC-Search Ainetter2024HOCSearch , ScanAnnotate ainetter2023automatically , Accurate Instance-Level CAD Model Retrieval Wei2022AccurateInstance , and FastCAD Langer2024FastCAD introduce similarity-based metrics that go beyond categorical correctness. Based on comparisons across these approaches, we adopt the L1 Chamfer Distance as our primary metric, measuring the geometric difference between the normalized retrieved CAD model and the normalized ground-truth model (as annotated in Scan2CAD). Our use of this metric is consistent with prior works Ainetter2024HOCSearch ; gao2024diffcadweaklysupervisedprobabilisticcad ; Wei2022AccurateInstance .
Unlike the Scan2CAD protocol, which limits retrieval to a predefined set of CAD candidates, we perform retrieval over the entire ShapeNet category corresponding to each ground-truth object. The number of CAD models per category used in our evaluation is listed in Table 4.
Table 4: Number of CAD model templates in the ShapeNet database for retrieval
Trash Can Bathtub Bookshelf Cabinet Chair Display Sofa Table
343 477 452 1571 6778 1093 3173 8436
Although some earlier works report similarity evaluations, direct comparison remains difficult due to missing details or unavailable implementations. To support reproducibility, we provide a comprehensive description of our evaluation methodology.
Representative Image Selection
Images of CAD models play a key role in retrieval pipelines that involve 2D feature extraction, silhouette matching, or depth comparisons. To select representative views for each CAD object, we reproject all 3D points of a scene instance into the image space and choose the top four images with the highest number of visible points—i.e., the least occluded views. For ShapeNet CAD models, we use the pre-rendered multi-view dataset from DISN Wang2019DISN , which contains 32 rendered images per model from different angles.
Chamfer Distance Computation
To compute Chamfer Distance, each CAD model is normalized—centered at the origin and scaled to fit within a unit cube (1$\times$1$\times$1). We uniformly sample 10,000 surface points from both the retrieved and ground-truth CAD models. The L1 Chamfer Distance is calculated bidirectionally as follows:
$\text{CD}_{L1}(A,B)=\frac{1}{|A|}\sum_{a\in A}\min_{b\in B}\|a-b\|_{1}+\frac{1}{|B|}\sum_{b\in B}\min_{a\in A}\|b-a\|_{1}$ (1)
The following three baselines are selected for comparison.
  • SCANnotate ainetter2023automatically : Utilizes an automatic CAD retrieval pipeline that leverages depth maps, silhouette comparison, and Chamfer Distance via differentiable rendering.
  • Digital Cousin dai2024acdc : Matches CAD models using 2D image features from multiple views to rank visually and semantically similar assets.
  • MSCD Wei2022AccurateInstance : Employs a two-step approach: initial retrieval using 3D feature descriptors, followed by geometry-based re-ranking using a Modified Single-direction Chamfer Distance (MSCD), which improves robustness to noise and better captures shape similarity in large databases.
We conduct the evaluation on all scenes from the ScanNetdai2017scannet validation set.
C.2 Object-Centric Material Recovery Benchmark
To evaluate object-centric material recovery in realistic conditions, we introduce a benchmark designed to reflect the challenges of real-world RGBD capture. Unlike global scene-level material estimation, our focus on object-level recovery aligns naturally with the object decomposition pipeline used in graphics, and enables per-object physically based rendering (PBR) material assignment—an essential representation for high-fidelity simulation and interactive graphics applications.
Figure 13: Challenging real-world conditions, including occlusion, poor lighting, skewed angles, and numerous distractors.
We constructed this benchmark using five indoor scenes captured with an iPhone 13 Pro Max via the 3D Scanner App in RoomPlan mode. These scenes cover a range of real-world environments, including a meeting room, a common area, a bedroom, and two study rooms, and contain a total of 111 manually curated objects spanning eight categories.
For each object, we identified and cropped the four most visually representative frames from the RGBD scan. These images often reflect real-world challenges such as occlusion, suboptimal lighting, and partial visibility—conditions under which we aim to test the robustness of various material recovery approaches. Some exampel crops are list in Figure 13. To benchmark performance, we compare our material painting pipeline against four baselines:
  1. 1.
    PhotoShape photoshape2018 – a material selection method using appearance-based retrieval;
  2. 2.
    Make-It-Real fang2024makeitreal – a language-guided material retrieval pipeline;
  3. 3.
    Our Visual+Language Search – combining visual similarity and textual prompts;
  4. 4.
    Make-It-Real + Albedo-Only Optimization – a variant incorporating our optimization stage.
Ground-truth CAD models are provided for all scenes. Estimated PBR materials from each method are applied to the CAD models and rendered at their original 3D poses using global illumination with HDR environment maps. For evaluation, we compare the rendered images of cropped objects against the corresponding cropped RGB images using perceptual metrics, as shown in Figure 14. To ensure fair evaluation, we manually selected retrieval results and object categories, avoiding retrieval-induced biases. Note that while we use known object poses for rendering and evaluation, these poses are not used during material recovery to preserve generalization to in-the-wild scenarios. We will release the scanned scenes, selected objects, cropped image sets, and evaluation scripts for reproducibility and support future research in object-centric material recovery.
Figure 14: Evaluation protocol used for assessing object-centric material painting performance.
C.3 Graphics-Ready Reconstruction Benchmark
We use five indoor scenes captured using an iPhone 13 Pro Max with the RoomPlan app, each containing hundreds of RGB frames. These scenes were also used in the Object-Centric Material Recovery benchmark. However, this benchmark evaluates a fully automated, end-to-end reconstruction pipeline—from raw input scans to a final graphics-ready scene. Lighting is not explicitly estimated; instead, we apply a general environmental light to provide consistent global illumination.
To measure visual quality, we adopt a perceptual similarity metric. Rendered images from the reconstructed scene—viewed under the original camera poses—are compared to real RGB captures.
We evaluate four reconstruction pipelines:
  1. 1.
    Phone2Proc phone2proc – a procedural reconstruction pipeline for embodied AI, operating at the room level with less emphasis on fidelity;
  2. 2.
    Digital Cousin dai2024acdc – integrated with our procedural reconstruction pipeline, but using the Digital Cousin flow for retrieval;
  3. 3.
    Digital Cousin + Make-It-Real (MIR) – enhanced with language-based material selection;
  4. 4.
    Digital Cousin + Ours (Vision+Semantic) – our proposed material selection method using combined vision-language guidance.
Appendix D More on Methodologies
D.1 Scene Parsing Stage
In this section, we describe the scene-parsing stage, which utilizes a scene graph to structure spatial relationships. The input includes detected wall structures and oriented object bounding boxes.
This is a rule-based approach, implemented in five key steps:
  • Step 1: Wall Closure. We start by taking the raw 2D wall segments and snapping any nearly matching endpoints together using a KD-Tree with a small distance threshold to close tiny gaps. We then build a simple graph in which each snapped point is a node and each wall segment is an edge. We traverse this graph to find a closed loop; if no loop exists, we connect the two loose ends and try again. The final loop becomes our room-boundary polygon.
  • Step 2: Wall–Object Alignment. Next, we examine each object’s 2D corners and orientation to determine which wall segment it belongs to. We select the closest wall (within 0.2 m) whose orientation is within 10° of the object’s facing direction. We then “snap” the object by moving its anchor point to the nearest point on that wall (leaving a small clearance) and rotating it to align its front face with the wall.
  • Step 3: In-Room Adjustment. To ensure no object sits outside the room, we test every object corner against the boundary polygon. Any object that falls outside triggers a local expansion of the offending wall segments and reconnection into a single loop. We repeat this until all objects are safely inside.
  • Step 4: Scene Graph Construction. At this stage, we compute wall connectivity and wall–object relationships. We then identify object–object relationships: if two objects’ projections onto the XY plane overlap and their Z extents indicate a clear top–bottom arrangement, we assign an “on-top” link; if two objects share the same orientation and lie within 0.1 m horizontally, we assign a “next-to” link.
  • Step 5: Collision Resolution. Finally, we resolve any remaining overlaps by treating each object footprint as a 2D polygon. Wall-attached objects are constrained to slide along their wall; others move freely. We iterate up to ten times, find every pair of overlapping objects, compute the minimal translation vector to push them apart (projecting along the wall if constrained), and move them until no collisions remain, or reach the maximum iteration.
We demensarate the results of scene parsing quantitiviely in Figure 17
D.2 Retrieval Stage
Figure 15: Detailed Breakdown of Procedural Generation The input data includes object nodes with spatial, aesthetic, and semantic information, which are utilized at different stages of the pipeline.
A detailed breakdown of each step in the proposed retrieval approach is illustrated in Figure 15.
Appendix E More Results
In this section, we present extended results of LiteReality, focusing particularly on the quantitative evaluation for some scenes in ScanNet, as shown in Figure 10.
For the individual stages of our pipeline, we first showcase the scene parsing results in Figure 17.
Regarding object retrieval, we provide a detailed analysis that includes both the reference images and the corresponding retrieved objects, as illustrated in Figure 18. We also highlight the effectiveness of our contextual selection module, and include some failure cases to demonstrate its limitations.
For the material painting component, we present further comparisons between our proposed method and several baseline models, with results visualized in Figure 16. Also the detailed per categories results tested on the proposed benchmark in Table 5.
Figure 16: Object-centric material painting results. Our method is compared against several baseline models, demonstrating improved texture alignment and realism.
Figure 17: Scene Parsing Enhancements in LiteReality. Visualization of several improvements introduced by our scene parsing module. Left and right panels show "Before" and "After" comparisons.
(a) Chair, Fireplace, Sofa, Storage
Method Chair Fireplace Sofa Storage
RMSE SSIM LPIPS RMSE SSIM LPIPS RMSE SSIM LPIPS RMSE SSIM LPIPS
PhotoShape 0.309 0.255 0.644 0.351 0.171 0.731 0.360 0.404 0.704 0.334 0.176 0.712
MIR+AO 0.227 0.403 0.582 0.285 0.403 0.677 0.270 0.489 0.639 0.218 0.389 0.609
MIR 0.242 0.384 0.604 0.291 0.394 0.668 0.263 0.488 0.645 0.238 0.365 0.629
Sem&Vis 0.263 0.372 0.627 0.299 0.449 0.665 0.266 0.475 0.685 0.287 0.339 0.652
LiteReality 0.228 0.407 0.578 0.283 0.433 0.688 0.270 0.486 0.634 0.230 0.395 0.621
(b) Table, Television, Door, Windows
Method Table Television Door Windows
RMSE SSIM LPIPS RMSE SSIM LPIPS RMSE SSIM LPIPS RMSE SSIM LPIPS
PhotoShape 0.401 0.304 0.640 0.351 0.242 0.652 0.259 0.311 0.658 0.441 0.215 0.750
MIR+AO 0.250 0.467 0.576 0.219 0.639 0.537 0.214 0.507 0.590 0.289 0.323 0.692
MIR 0.256 0.448 0.600 0.218 0.632 0.543 0.207 0.503 0.603 0.332 0.285 0.689
Sem&Vis 0.344 0.391 0.626 0.261 0.613 0.618 0.234 0.519 0.619 0.482 0.193 0.747
LiteReality 0.248 0.477 0.573 0.220 0.636 0.531 0.212 0.546 0.581 0.286 0.361 0.689
Table 5: Per-category evaluation metrics (RMSE ↓, SSIM ↑, LPIPS ↓) for different methods across various object types.
Figure 18: Demonstration of retrieval results using the proposed multi-step object retrieval pipeline. The image on the left displays the top four visible images from the dataset. The middle images show the top four ranked results from the visual selection process after Step 3 or Step 2. The right images present the results proposed by GPT-4o. GPT-4o leverages more contextual information, often making more reasonable selections compared to purely visual-based methods. However, failure cases can still occur, as shown in the bottom image, where unreasonable 3D model selection hinders the realism of LiteReality. With advancements in reasoning within MLLMs, we anticipate further improvements in such pipelines in performance as MLLMs continue to evolve.
Appendix F Application of LiteReality
In this section, we showcase how scenes reconstructed with LiteReatliy can be used for a wide range of application, benifits from its graphic realted featuers. The Demensration includes:
Relighting: Reconstructed scenes with physically based rendering (PBR) materials support dynamic relighting under various conditions, as demonstrated in Figure 19. This adaptability enhances rendering fidelity and enables realistic, immersive experiences.
Figure 19: Lighting Effects on Procedural Reconstructions: This figure illustrates the versatility of procedural reconstructions enhanced with PBR materials under various lighting conditions. The top row showcases the reconstructed scene rendered under different lighting setups: standard indoor lighting, TV lighting where the television acts as the sole light source, morning lighting with natural daylight streaming through the window, and dawn lighting that mimics the soft glow of early sunrise. The bottom row provides a comparison to the original RGBD scans used as inputs. This comparison demonstrates the effectiveness of the pipeline in achieving a realistic scene appearance. The ability to seamlessly re-light scenes post-reconstruction highlights the utility of this approach in diverse applications, such as virtual reality, architectural visualization, and cinematic production, where dynamic lighting enhances realism and usability. [Scene is reconstructed from ScanNet scene0166]
Physics-Based Interactions: The pipeline enables realistic rigid-body physics and collision detection, allowing natural interactions such as objects falling or colliding. Figure 20 illustrates these capabilities with examples like a falling apple and dynamic indoor collisions.
Figure 20: Physics-Based Interaction in Procedural Reconstructions: This figure demonstrates the application of physics simulations to procedurally reconstructed scenes, enabling realistic interactions between objects. The left panel highlights two examples: (1) a gravity simulation where numerous apples are dropped into the scene, naturally landing on tables, scattering, and interacting with the environment; and (2) a toy car driving at high speed, colliding with furniture, showcasing rigid body dynamics as individual objects respond realistically to impacts and forces. The right panel provides a comparison of the original RGB images, reconstructed geometry, and final simulation-ready scenes, emphasizing how the reconstructed objects exhibit realistic physical behaviour. This integration of physics-based interactions enhances the realism and utility of reconstructed scenes, making them suitable for applications in simulation environments, gaming, and robotics.[Scene is reconstructed from ScanNet scene0291]
Scene Editing: Modular reconstruction allows users to dynamically replace, modify, or add objects. Integrated physics ensures realistic placement and interaction, simplifying tasks such as adding messy, lifelike elements to a scene (Figure 21).
Applications in Robotics, Gaming, and VR: The structured and interactive nature of reconstructed scenes facilitates seamless integration into robotics, gaming, and VR environments. Figure 21 showcases examples that demonstrate the pipeline’s potential for creating realistic digital replicas of real-world spaces.
Figure 21: Demonstration of applications enabled by graphics-ready reconstructed scenes, showcasing advanced features such as physics-based interactions, robotic training, gaming environments, and immersive virtual reality experiences. These examples highlight the versatility and interactivity of the reconstructed scenes for various use cases. (Scene is reconstructed from ScanNet Scene0281)
Appendix G Limitations and Future Work
Limitations
While LiteReality converts RGB–D scans into compact, interactive 3D scenes, it still has several limitations. First, our procedural reconstruction relies on hand-tuned rules and artist-crafted assets, which can fail in unconventional or highly customized layouts (e.g., a sink flush-mounted within bespoke cabinetry). Second, although our Material Painting module recovers high-quality, spatially varying albedo maps, it does not estimate illumination or specular properties, limiting relighting fidelity under novel lighting conditions. Third, small or cluttered objects—such as mugs, books, and cables—are currently omitted, reducing scene completeness. Fourth, reconstruction quality remains tightly coupled to upstream perception: errors in object detection can propagate. Finally, our simple constraint-based collision solver can struggle in complex arrangements, occasionally producing interpenetrations and preventing fully physics-valid simulations.
Future Work
Integrating small-object detection and retrieval into the pipeline would enrich scene detail and realism. Incorporating learning-based light estimation together with full SVBRDF recovery would enable faithful relighting under novel illuminations. Exploring learned or physics-driven collision-resolution techniques could ensure non-overlapping, simulatable layouts. Finally, scaling the approach from individual rooms to multi-room and building-scale reconstructions would unlock applications in large-scale visualization and digital-twin creation.